# Deep Learning for Pathological Speech: A Survey

### v1

@@ IX-B Graph Neural Networks for Pathological Speech Detection
While many automatic pathological speech detection methods have demonstrated remarkable performance, the exploitation of common attributes such as pathological cues, speaker characteristics, age, and gender across different speakers remains limited. A promising direction would be to approach the pathological speech detection problem as a semi-supervised node graph classification task via graph neural networks (GNNs). This could involve constructing an inter-speaker graph from the utterances of various speakers, where the graph connectivity would help form speaker clusters based on the presence or absence of pathological cues. This structural information can then be leveraged through various graph message-passing mechanisms to improve detection accuracy and provide deeper insights into pathological speech patterns. Initial work in this direction has been done in [218], where a graph convolutional neural networks has been proposed for Parkinson’s disease detection method by framing the task as a node classification problem. In this approach, a graph is constructed using w2v2 speech embeddings across different speech segments from various patients as nodes, with the edges capturing the relationships between these segments. We believe that in this context, a promising research direction is to explore the incorporation of domain knowledge, such as establishing edges based on factors like gender, age, and severity scale of the patients.
IX-C Privacy
@@
In pathological speech-based applications such as detection models or ASR systems, privacy-preserving solutions are crucial for safeguarding sensitive patient data. Since these systems process speech that reveals sensitive medical information, ensuring data security and confidentiality is of paramount importance. By employing privacy-preserving techniques - such as federated learning, differential privacy, and encryption - providers can ensure that individuals’ speech data is anonymized and never exposed to unauthorized parties. This not only fosters trust among patients but also complies with stringent data protection regulations, thereby mitigating the risk of breaches. Developing privacy-preserving pathological speech processing systems where one must balance two conflicting goals, i.e., increasing the utility of the models while preserving the privacy of the users, remains an important challenging research directions.
IX-D Multimodal Pathological Speech Analysis
@@
Most existing systems primarily focus on leveraging speech cues for detecting pathological speech. However, complementary cues may also exist in visual forms, such as facial expressions or lip movements, which can provide additional insights. To effectively utilize these visual cues, multimodal frameworks like AV-HuBERT could prove advantageous, as they integrate both audio and visual information, potentially enhancing the accuracy and robustness of pathological speech detection models. Furthermore, combining other modalities, such as electroencephalogram (EEG) brain signals and textual data (transcripts), could offer a more complete characterization of disorders. These multimodal systems may help in capturing underlying neural, visual or linguistic patterns associated with speech pathology, improving detection and diagnosis across diverse patient groups.
X Explainability and Interpretibility
In the domain of pathological speech detection, explainability and interpretability are crucial due to their clinical significance. Although these terms are often used interchangeably because of their close overlap, it’s important to distinguish between them. As defined by [219], interpretable models are designed to be inherently understandable, whereas explainable models provide post-hoc explanations for the decisions made by existing ”black box” systems that are otherwise incomprehensible to humans. [220] further clarified the distinction by defining interpretability as ”How does the model work” and explainability as ”What else can the model tell me.” Keeping these distinctions in mind, particularly in the context of pathological speech detection, there has been very limited, if any, attention given to this aspect of research [221, 222, 223, 224]. Most existing studies focus on improving model accuracy and performance, often overlooking the importance of making these models interpretable and explainable. This gap is critical, as the ability to understand and explain the decision-making process in pathological speech detection models can provide valuable insights for clinicians and support more trustworthy and reliable use of artificial intelligence in healthcare applications. To address this issue, [221] attempted to map high-dimensional acoustic features to sound pattern of English binary phonological features, recognizing that these high-dimensional acoustic features are challenging to interpret in clinical practice. The problem with this approach is that it may be biased toward English sounds, making it susceptible to errors when applied to different phonetic environments. Further, [224] applied the SHAP algorithm to identify the most influential features in their model and discovered that constant-vowel transitions were the key features in reversing the model’s decisions in pathological speech detection. Recently, [225] demonstrated using canonical correlation analysis that the human frequency range (0–210 Hz) plays a critical role in influencing model decisions for pathological speech detection. Although some research has been conducted, substantially more research is needed to address this under-explored but very important area.
XI Large Language Models
for Pathological Speech
@@ XI Large Language Models
Given the recent advancements in large language models (LLMs), which have demonstrated significant progress across a wide range of applications [226, 227], exploring their potential for pathological speech analysis appears to be an inevitable and promising direction. LLMs could be leveraged for tasks such as detecting speech impairments based on specific phrase cues, improving the automatic classification of speech disorders, and enhancing the reconstruction of unintelligible or difficult speech. Additionally, LLMs’ ability to capture complex linguistic patterns may lead to more accurate models for personalized therapy, rehabilitation, and assistive communication tools for individuals with speech impairments. This exploration could bridge the gap between traditional speech processing techniques and state-of-the-art language models, opening up new avenues for more effective and adaptive speech rehabilitation systems.
XII Conclusion
@@
This paper provides a comprehensive overview of deep learning applications for pathological speech, encompassing detection, recognition, intelligibility assessment and enhancement. Additionally, it compiles a thorough list of both accessible and non accessible pathological speech datasets, which will serve as valuable resources for future research and accelerate progress in the field. To the best of our knowledge, this is the first review to offer an in-depth survey of pathological speech from the perspectives of detection, recognition, enhancement, explainibility and data augmentation. Furthermore, it outlines potential future research directions, particularly in the context of robust and interpretable models that can be deployed in clinical practice.

### v2

@@ IX-C Improving the Performance of Automatic Pathological Speech Analysis
While many automatic pathological speech detection methods have shown remarkable performance, the exploitation of common attributes such as pathological cues, speaker characteristics, age, and sex across different speakers remains limited. A promising direction is to approach the pathological speech detection problem as a semi-supervised node graph classification task using graph neural networks (GNNs), as demonstrated in [231]. This approach could involve constructing an inter-speaker graph based on utterances from various speakers, where the graph’s connectivity would help form speaker clusters based on the presence or absence of pathological cues. In this context, domain knowledge can be easily integrated by establishing edges based on factors such as sex, age, and the severity scale of the patients. Additionally, the recent availability of longitudinal multimodal data [61] enables novel applications of GNNs in disease monitoring. By representing the patient history as a temporal graph where nodes capture multimodal features (e.g., vocal biomarkers, facial expressivity) and edges encode their dynamic interactions, GNNs can model disease progression through evolving graph topologies.
An additional area that remains under-explored in current research, potentially due to the lack of large datasets, is the adoption of Bayesian frameworks and generative approaches for pathological speech processing. Bayesian frameworks are particularly relevant in healthcare, as they allow the model to estimate uncertainties in its predictions, which is essential for high-risk settings where incorrect predictions can have serious consequences. These models are ideal for situations where speech data may vary considerably due to speaker differences or environmental factors. By providing probabilistic models, Bayesian methods can account for such variability, making them highly suitable for clinical applications where individual differences are pronounced.
IX-D Privacy
@@
In pathological speech-based applications such as detection models or ASR systems, privacy-preserving solutions are crucial for safeguarding sensitive patient data. Since these systems process speech that reveals sensitive medical information, ensuring data security and confidentiality is of paramount importance. By employing privacy-preserving techniques such as federated learning, differential privacy, and encryption, providers can ensure that individuals’ speech data is anonymized and never exposed to unauthorized parties. This not only fosters trust among patients but also complies with stringent data protection regulations, thereby mitigating the risk of breaches. Developing privacy-preserving pathological speech processing systems where one must balance two conflicting goals, i.e., increasing the utility of the models while preserving the privacy of the users, remains an important challenging research directions.
IX-E Multimodal Pathological Speech Analysis
@@
Most existing systems primarily focus on leveraging speech cues for detecting pathological speech. However, complementary cues may also exist in visual forms, such as facial expressions or lip movements, which can provide additional insights. To effectively utilize these visual cues, multimodal self-supervised methods such as AV-HuBERT [237] could prove advantageous, as they facilitate the integration of multimodal audio and visual information, potentially enhancing the accuracy and robustness of pathological speech detection models. Furthermore, combining other modalities, such as medical imaging, health records, brain signals, and textual data, could offer a more complete characterization of disorders. These multimodal systems may help in capturing underlying neural, visual or linguistic patterns associated with speech pathology, improving pathological speech processing across diverse patient groups.
IX-F Explainability and Interpretability
In the domain of pathological speech detection, explainability and interpretability are crucial due to their clinical significance. Although these terms are often used interchangeably, it is important to distinguish between them. As defined in [238], interpretable models are designed to be inherently understandable, whereas explainable models provide post-hoc explanations for the decisions made by existing black box systems that are otherwise incomprehensible to humans.
Despite their importance, relatively little attention has been paid to explainability and interpretability in pathological speech detection, with most existing research focusing primarily on improving model accuracy. Nonetheless, a few notable efforts have begun to address this gap [239, 240, 241, 242, 243]. For example, [239] mapped high-dimensional acoustic features to binary phonological representations, recognizing that raw acoustic features are difficult to interpret in clinical practice. Xu et al. [243] applied the SHAP algorithm to identify the most influential features in their model, highlighting the role of consonant-vowel transitions in reversing classification decisions. Similarly, [232] used canonical correlation analysis to show that the 0210210 Hz frequency range strongly influences model outputs.
Recently, [151] found that non-interpretable SSL embeddings outperform interpretable features (e.g., prosodic, linguistic, and cognitive descriptors) in both multilingual and cross-lingual contexts. This highlights a growing challenge, i.e., as models increasingly rely on high-performing but opaque representations like SSL embeddings, there is a pressing need to develop methods that make these models explainable and clinically trustworthy.
IX-G Large Language Models for Pathological Speech
@@
Given the recent advancements in multimodal large language models (LLMs), which have demonstrated significant progress across a wide range of applications [244, 245], exploring their potential for pathological speech analysis appears to be an inevitable and promising direction. Multimodal LLMs could be leveraged for tasks such as detecting speech impairments, ASR, and enhancing the reconstruction of unintelligible or difficult speech. Their ability to capture complex patterns may lead to more accurate models for personalized therapy, rehabilitation, and assistive communication tools for individuals with speech impairments. Additionally, multimodal LLMs should be explored for their ability to explain their decisions. This exploration could bridge the gap between traditional speech processing techniques and state-of-the-art language models, opening up new avenues for more effective and adaptive speech rehabilitation systems.
X Conclusion
@@
This paper provides a comprehensive overview of speech analysis and technologies for pathological speech arising due to neurological disorders, encompassing detection, recognition, intelligibility assessment, and enhancement. Additionally, it compiles a thorough list of both accessible and non-accessible pathological speech datasets, which will serve as valuable resources for future research and accelerate progress in the field. Finally, it outlines potential future research directions, particularly in the context of robust and interpretable models that can be deployed in clinical practice.